A^{2}Π ˜ X^{2}Σ^{+} INTERACTION PARAMETERS FOR ^{12}C^{16}O^{+}, ^{13}C^{16}O^{+} AND ^{14}C^{16}O^{+} FROM DEPERTUR-BATION ANALYSES OF A^{2}Π(v′) → X^{2}Σ^{+}(v′′) BANDS

\underline{JOHN A. COXON}, Department of Chemistry, Dalhousie University, Halifax, Nova Scotia B3H 4J3, Canada; R. KȨPA, A. KOCAN and I. PIOTROWSKA-DOMAGALA, Atomic and Molecular Physics Laboratory, Institute of Physics, Univerity of Rzeszów, 35-310 Rzeszów, Poland.

Several A^{2}Π → X^{2}Σ^{+} emission bands of CO^{+}, for which the A-state vibrational level is perturbed, have been recorded photographically. The well known A(υ′ = 0) ~ X(υ^{∗} = 10) and A(υ′ = 5) ~ X(υ^{∗} = 14) perturbations between the A^{2}Π and X^{2}Σ^{+} states of ^{12}C^{16}O^{+}, previously analysed by Coxon and Foster^{a}, have been reinvestigated by analysis of the 0 - 2 and 5 - 0 bands. The standard deviations of the least squares fits are typically about 0.02 cm^{−1}. Similarly, the A(υ′ = 1) ~ X(υ^{∗} = 11) interaction in ^{13}C^{16}O^{+}, first identified by Jakubek^{b} in an analysis of the 0 - 1 band of the B^{2}Σ^{+} − A^{2}Π system, has been studied by a deperturbation analysis of the rotational structure in the A^{2}Π(υ′ = 1) → X^{2}Σ^{+}(υ′′ = 0) band. For ^{14}C^{16}O^{+}, the corresponding A(υ′ = 2) ~ X(υ^{∗} = 11) interaction has been observed for the first time by analysis of the 2 - 0 and 2 - 1 bands of the A − X system. The expected isotopic self-consistency of the interaction parameters α and β from individual bands of the three isotopomers is discussed.

aJ. A. Coxon and S. C. Foster, J. Mol. Spectrosc. 93, 117 (1982).

bZ. Jakubek, J. Mol. Spectrosc. 131, 207 (1988).